A manufacturer who produces medicines bottles finds that 0.1% of the bottles are defective. The bottles are packed in boxes containing 500 bottles. A drug manufacturer buys 100 boxes from producer of bottles. Using possion distribution, how many bottles will contain

(1) no defectives (2) at least two defectives


1
Expert's answer
2021-02-17T13:17:50-0500

We have:

"N=100, n=500"

Probability of a defective bottle:

"p=0.1\%=0.001"

"\lambda=np=500\cdot0.001=0.5"

If "X" denotes the number of defective bottles in a pack of 500, then by Poisson distribution:

"P(X=x)=e^{-0.5}\frac{0.5^x}{x!}=\frac{0.6065\cdot0.5^x}{x!}" ; "x=0,1,2,..."

Hence in a lot of 100 boxes the frequency of boxes with "x" defective bottles is given by

"f(x)=NP(X=x)=\frac{100\cdot0.6065\cdot0.5^x}{x!}"


1) Number of bottles with no defectives:

"=100P(X=0)=100\cdot0.6065=61"


2) Number of bottles with at least two defectives:

"=100P(X\geq2)=100(1-P(X=0)-P(X=1))="

"=100(1-0.6065-0.6065\cdot0.5)=9"